E.12.3 investment in an item of equipment is $22,000. It has a five-year life and no salvage value and shirt one depreciation method is used the equipment is expected to provide an annual savings of $2900, which does not include depreciation. What is the pay back period?
Answer #1

Cost of equipment = $22,000

Useful life = 5 years

Annual depreciation = Cost of equipment / Useful life

= 22,000/5

= $4,400

Annual savings = $2,900

Annual cash inflow = Annual depreciation+Annual savings

= 4,400+2,900

= $7,300

Pay back period = Cost of equipment / Annual cash inflow

= 22,000/7,300

= 3.01 years or 3 years (if rounded to nearest whole year)

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks